OpenStudy (anonymous):

2yx=7y+5 with point (-9,2) what does y equal in slope intercept form

OpenStudy (anonymous):

slope intercept form: y = mx + b 2yx-7y = 5 y(2x-7)=5 y=5/(2x-7) Is that complete enough? I suppose you could put in (-9,2) to solve for B, but i'm not sure if that's necessary considering the initial +5 constant given.

OpenStudy (anonymous):

is that 5 divided by (2x-7)

OpenStudy (anonymous):

Yes.

OpenStudy (vijay):

nope, he need y=mx+c format

OpenStudy (anonymous):

I need to the equation in slope intercept form

OpenStudy (anonymous):

Right, well i'm not sure how to get that from 5/(2x-7). I'll think about it.

OpenStudy (anonymous):

thanks @vijay any help